John Scott has R200 per week available to spend on beer and cigarettes. Beer costs R10 per bottle and cigarettes cost R40 per packet. Suppose John buys 8 bottles of beer and 3 packets of cigarettes and that his marginal utility from beer is then 20 utils and that from cigarettes 35 utils. Is he in equilibrium? If not, should he buy more beer or more cigarettes? Explain.
1
Expert's answer
2019-04-03T09:04:45-0400

In equilibrium:

"\frac{MU1} {P1} = \frac{MU2} {P2}"

and

"P1 \times Q1 + P2 \times Q2 = I"

In our case 20/10>35/40, so John should consume more beer and less cigarettes to maximize his utility.